# Seismocardiography for Emotion Recognition: A Study on EmoWear with Insights from DEAP

### v2

@@ V-B Interpretation of Findings
The most interesting insights are observed when the SCG signal is combined with the ADR signal, as this combination uses only a single chest-worn accelerometer to extract both signals. The boxplots of the ‘SCG + ADR’ combination in Figure 7 show an evident drop of performance comparing to their adjacent boxplots, meaning that the ADR signal has not been able to reach the same level of performance as the RSP / all signals. Figure 8 provides a more detailed view on the cardio-respiratory-based combination results. Figure 8a confirms that the average $F1$ scores of the ‘SCG + ADR’ combination is comparable to the ‘ECG / BVP + RSP’ combinations, but slightly lower. Figure 8b shows the Pearson correlation heatmap of the $F1$ scores obtained across the four cardio-respiratory signal combinations. The heatmap suggests that the ‘SCG + ADR’ combination performs differently compared to the other signal combinations, within subjects. Noteworthily, the different distribution of the ‘SCG + ADR’ combination, does not mean that the combination is not effective. The t-test results in Table VI reveal otherwise: while the NB classifier fails to achieve significant results, the more robust SVM and LR classifiers demonstrate $F1$ score distributions that are significantly higher than best of the baselines ($p<0.001$; see Table VI, rows 20–21). These findings confirm the applicability of a single chest-worn accelerometer for ER, which is a significant step towards embedding ER in everyday life scenarios.
@@ V-D Future Directions
Outcomes and limitations seen in this study suggest several potential directions for future research. These directions include but are not limited to: 1) Exploring the potential of the SCG signal for ER in more detail, including the development of dedicated feature extraction methods and models; 2) Investigating and improving signal processing techniques for more robust information inference from the SCG and ADR signals; 3) Exploring deep learning approaches like CNNs, LSTMs, or transformers to process raw SCG signals directly, circumventing limitations of manual feature engineering; 4) Exploring transfer learning or self-supervised pre-training on large-scale physiological datasets to overcome the challenge of limited dataset size; 5) Investigating subject-independent ER problem using stronger models, capable of learning more complex patterns; 6) Uncovering the potential of chest-worn accelerometers for ER in everyday life scenarios, including the development of real-time ER systems; and 7) Exploring the potential of the EmoWear dataset for other applications, such as voice activity detection, drinking detection, and activity recognition, which are also included in the dataset.

### v3

@@ III-A Dataset
Subject Demographics
The DEAP dataset contains 32 subjects (15 females and 17 males) aged between 19 and 37 years (mean=26.9), while the EmoWear dataset contains 48 subjects (21 females and 27 males) aged between 21 and 45 years (mean=29.3). One subject in the DEAP dataset reported a history of migraine, while three subjects in the EmoWear dataset reported histories of ADHD - autism, panic attacks, and epilepsy.
@@ V-A Methodological Approach
When comparing with existing accelerometer emotion recognition research, we note that this study is the first to analyze accelerometer data as a mechanical surrogate of cardiac activity (SCG) for ER. The use of accelerometer data for ER was previously limited to other motion characterization methods, such as blind motion measurement [80, 81, 82] or gate-analysis [35]. Using SCG provides a more direct measurement of the heart’s mechanical activity, which is less affected by motion artifacts than other accelerometer-based methods.
@@ V-B Interpretation of Findings
The most interesting insights are observed when the SCG signal is combined with the ADR signal, as this combination uses only a single chest-worn accelerometer to extract both signals. The boxplots of the ‘SCG + ADR’ combination in Figure 7 show an evident drop of performance comparing to their adjacent boxplots, meaning that the ADR signal has not been able to reach the same level of performance as the RSP / all signals. Figure 8 provides a more detailed view on the cardio-respiratory-based combination results. Figure 8a confirms that the average $F1$ scores of the ‘SCG + ADR’ combination is comparable to the ‘ECG / BVP + RSP’ combinations, but slightly lower. We do not see this performance drop when the SCG is combined with the RSP signal. Therefore, the slight drop in performance of the ‘SCG + ADR’ combination can be attributed to the fact that the ADR signal tends to be noisier and less specific than a dedicated respiratory belt (RSP), limiting the signal fidelity. Figure 8b shows the Pearson correlation heatmap of the $F1$ scores obtained across the four cardio-respiratory signal combinations. The heatmap suggests that the ‘SCG + ADR’ combination performs differently compared to the other signal combinations, within subjects. Noteworthily, the different distribution of the ‘SCG + ADR’ combination, does not mean that the combination is not effective. The t-test results in Table VI reveal otherwise: while the NB classifier fails to achieve significant results, the more robust SVM and LR classifiers demonstrate $F1$ score distributions that are significantly higher than best of the baselines ($p<0.001$; see Table VI, rows 20–21). These findings confirm the applicability of a single chest-worn accelerometer for ER, which is a significant step towards embedding ER in everyday life scenarios.
@@ V-D Future Directions
Outcomes and limitations seen in this study suggest several potential directions for future research. These directions include but are not limited to: 1) Exploring the potential of the SCG signal for ER in more detail, including the development of dedicated feature extraction methods and models; 2) Investigating and improving signal processing techniques for more robust information inference from the SCG and ADR signals; 3) Examining the potential of lightweight deep learning models (e.g., MobileNet, TinyLSTM) for ER on the EmoWear dataset, in addition to the explored traditional classifiers; 4) Exploring deep learning approaches like CNNs, LSTMs, or transformers to process raw SCG signals directly, circumventing limitations of manual feature engineering; 5) Exploring transfer learning or self-supervised pre-training on large-scale physiological datasets to overcome the challenge of limited dataset size; 6) Investigating subject-independent ER problem using stronger models, capable of learning more complex patterns; 7) Uncovering the potential of chest-worn accelerometers for ER in everyday life scenarios, including the development of real-time ER systems; 8) Exploring the potential of the EmoWear dataset for other applications, such as voice activity detection, drinking detection, and activity recognition, which are also included in the dataset; 9) Expanding the existing dataset by collecting more data from a larger and more diverse group of participants, to improve the generalizability of the findings; and 10) Developing a chest-worn smartphone application that can be used by participants in their daily lives, to collect more data in an opportunistic manner.